# Exploring neural oscillations during speech perception via surrogate gradient spiking neural networks

### v1

@@ Results
In order to draw a comparison with the human auditory pathway, we introduce a physiologically inspired ASR pipeline illustrated in Fig. 1. The proposed hybrid ANN-SNN architecture is trained in an end-to-end fashion on the TIMIT dataset [50] to predict phoneme sequences from speech waveforms. In the architectural design, we aimed to minimise the complexity of ANN components and favour the central SNN which will be the focus of the oscillation analysis. Here as a preliminary step, we examine how relevant hyperparameters affect the phoneme error rate (PER). On top of assessing the scalability of our approach to larger networks, we identify the importance of the interplay between recurrence and SFA.
Network scalability
@@
As reported in Table 1, performance improves with added layers, peaking at 4-6 layers before declining, which suggests a significant contribution to the final representation from all layers within this range. Compared to conventional non-spiking recurrent neural network (RNN) encoders used in ASR, our results support the scalability of surrogate gradient SNNs to relatively deep architectures. Additionally, augmenting the number of neurons until about 1,000 per layer consistently yields lower PERs, beyond which performance saturates.
@@
Table 1: Hyperparameter tuning for the number of SNN layers and neurons per layer on the TIMIT dataset. The second column gives both the number of trainable parameters in the multi-layered SNN (left) and in the whole encoder (right). The PERs are reported after 50 training epochs using a 5 ms time step, 16 convolutional neural network (CNN) channels, 50% of adaptive leaky integrate-and-fire (AdLIF) neurons, 100% feedforward and 50% recurrent connectivity. The performance of the architecture when removing the SNN is also reported (bottom). Bold values indicate the lowest achieved PERs.
@@
Recurrent connections and spike-frequency adaptation
@@
Corrolary
@@
We observe in Supplementary Table 1 that decreasing the simulation time step does not affect the performance. Although making the simulation of spiking dynamics more realistic, one might have anticipated that backpropagating through more time steps could hinder the training and worsen the performance as observed in standard RNNs often suffering from vanishing or exploding gradients [52]. With inputs ranging from 1,000 to over 7,000 steps using 1 ms intervals on TIMIT, our results demonstrate a promising scalability of surrogate gradient SNNs for processing longer sequences. Secondly, as reported in Supplementary Table 2, we did not observe substantial improvement when increasing the number of auditory nerve fibers past $\sim$5,000, even though there are approximately 30,000 of them in the human auditory system. This could be due to both the absence of a proper model for cochlear and hair cell processing in our pipeline and to the relatively low number of neurons (<1,000) in the subsequent layer. Finally, as detailed in Supplementary Table 3, reduced feedforward connectivity in the SNN led to poorer overall performance. This contrasts with our earlier findings on recurrent connectivity, highlighing the distinct functional roles of feedforward and feedback mechanisms in the network.
Oscillation analysis
@@
Based on our previous architectural results that achieved satisfactory speech recognition performance using a physiologically inspired model, we hypothesise that the spiking dynamics of a trained network should, to some extent, replicate those occurring throughout the auditory pathway. Our investigation aims to discern if synchronisation phenomena resembling brain rhythms manifest within the implemented SNN as it processes speech utterances to recognise phonemes.
Synchronised gamma activity produces low-frequency rhythms
@@
As illustrated in Fig. 22, the spike trains produced by passing a test-set speech utterance through the trained architecture exhibit distinct low-frequency rhythmic features in all layers. By looking at the histogram of single neuron firing rates illustrated in Fig. 22, we observe that the distribution predominantly peaks at gamma range, with little to no activity below beta. This reveals that the low-frequency oscillations visible in Fig. 22 actually emerge from the synchronisation of gamma-active neurons, which is understood to be a key operational feature of activated cortical networks. The resulting low-frequency rhythms appear to follow to some degree the intonation and syllabic contours of the input filterbank features and to persist across layers. Compared to the three subsequent layers, higher activity in the auditory nerve comes from the absence of inhibitory SFA and recurrence mechanisms. These initial observations suggest that the representation of higher-order features in the last layer is temporally modulated by lower level features already encoded in the auditory nerve fibers, even though each layer is seen to exhibit distinct rhythmic patterns. In the next section, we focus on measuring this modulation more rigorously via PAC analysis.
@@
[…]
Figure 2: Spiking activity in response to speech input. a Input filterbank features and resulting spike trains produced across layers. For each layer, the neurons are vertically sorted on the y-axis by increasing average firing rate (top to bottom). The model uses a 2 ms time step, 16 CNN channels, 3 layers of size 512, 50% AdLIF neurons, 100% feedforward and 50% recurrent connectivity. b Corresponding distribution of single neuron firing rates.
Phase-amplitude coupling within and across layers
By aggregating over the relevant spike trains, we compute distinct EEG-like population signals for the auditory nerve fibers and each of the three SNN layers. These are then filtered in the different frequency bands, as illustrated in Fig. 33, which allows us to perform CFC analyses. We measure PAC within-layer and across-layers between all possible combinations of frequency bands and find multiple significant forms of coupling for every utterance. An example of significant theta low-gamma coupling between the auditory nerve fibers and the last layer is illustrated in Fig. 33. Here input low-frequency modulation is observed to significantly modulate the output gamma activity. The network therefore integrates and propagates intonation and syllabic contours across layers via a synchronised neural activity along these perceptual cues. It is important to note that the synchronisation of neural signals to the auditory envelope emerged without imposing any theta or gamma activity in our network. The optimisation of the PER combined with sufficiently realistic spiking neuronal dynamics therefore represent sufficient conditions to reproduce some broad properties of neural oscillations observed in the brain, suggesting a general functional role of facilitating information processing. More generally, the patterns of coupling between different neural populations and frequency bands were found to differ from one utterance to another. These variations indicate that the neural processing of our network is highly dynamic and depends on the acoustic properties and linguistic content of the input. The observed rich range of intra- and inter-layer couplings suggests that the propagation of low-level input features such as intonation and syllabic rhythm is only one part of these synchronised neural dynamics.
[…]
Figure 3: Cross-frequency coupling of population aggregated activity. a Population signals of auditory nerve fibers (blue) and last layer (orange) filtered in distinct frequency bands. b Modulation index and mean vector length metrics as measures of PAC between the theta band of the auditory nerve fibers and the low-gamma band of the last layer.
Impact of spike-frequency adaptation and recurrent connections on oscillations
As presented in Table 3, we record an overall average activity of 100 Hz in a trained architecture with no SFA and no recurrent connections. Using recurrent connections but no SFA, the average activity drops to 71 Hz where 60% of trained recurrent weights are inhibitory (see Supplementary Fig. 2) and help regulate the activity. Using SFA with or without recurrent connections further decreases the firing rate to 61 Hz emphasising the overall inhibitory effect of SFA. This regularisation of the network activity is able to improve the encoding of speech through the architecture as it results in better ASR performance. While both forms of feedback exhibit an overall inhibitory effect, SFA operates at the individual neuron level whereas recurrent connections act at the layer level. Our analysis indicates distinct influences of these two mechanisms on CFCs. On the one hand, inhibitory feedback produced by SFA reinforces synchronisation rhythms in both intra- and inter-layer interactions, especially within theta and alpha bands. This is in line with the work of Crook et al. [53], which demonstrated that SFA can encourage and stabilise the synchronisation of cortical networks. Similarly, Augustin et al. [54] showed that adaptation promotes periodic signal propagation. In our case, the resulting temporal reorganisation and activity regulation occurring at the neuron level appear to enable more effective parsing and encoding of speech information, leading to improved ASR performance. It is interesting to note that the adaptation time constant $\tau_{w}\in[30,350]$ ms corresponds to delta-beta ranges, whereas the membrane potential time constant $\tau_{u}\in[3,25]$ ms corresponds to gamma range. SFA therefore naturally enables the production of gamma bursts occurring at lower frequencies. On the other hand, networks with recurrent connections are associated with an increase in intra-layer couplings especially in the last layer, yet a decrease in inter-layer couplings. Layer-wise recurrence therefore enables more complex intra-layer dynamics but reduces inter-layer synchronisation, resulting in more localised rhythms.
Table 3: Comparison of the oscillatory activity resulting from passing the 64 longest TIMIT test-set utterances through networks with and without SFA and recurrent connections. The first column reports the average firing rate of each network. Columns delta to beta then give the number of significant PACs measured for each modulating frequency band across all intra- and inter-layer interactions. The last column gives the total number of significant PACs across all frequency bands. All networks use a 2 ms time step, 16 CNN channels, 3 layers, 512 neurons per layer and 100% feedforward connectivity.
Model Type
Firing rate
a[Hz]
Delta
(intra, inter)
Theta
(intra, inter)
Alpha
(intra, inter)
(intra, inter)
Total
(intra, inter)
No recurrence no SFA 100 (11, 21) (39, 73) (61, 69) (22, 29) (133, 192)
Recurrence only 71 (12, 3) (67, 39) (62, 61) (39, 29) (180, 132)
SFA only 61 (14, 29) (53, 161) (102, 162) (44, 38) (213, 390)
Recurrence and SFA 61 (12, 13) (64, 51) (103, 79) (86, 34) (265, 177)
Table 4 outlines the five most prominent coupling patterns observed within each architecture. The activity of the last SNN layer stands out as the most significantly modulated overall. By architectural construction, modulation in that final layer has the most impact on the ASR task as the corresponding spike trains are the ones converted to phoneme probabilities using a small ANN. In the last layer, an increase in the occurrences of theta, alpha and beta intra-layer couplings goes together with a decrease of the PER, which suggests more selective integration of phonetic features, enhanced attentional processes, as well as improved assimilation of contextual information.
Table 4: Most measured coupling types. After analysing the 64 longest utterances of TIMIT test-set, we report the five types of coupling for which significant PAC was measured on most utterances. Models with and without SFA and recurrent connections are compared.
Model type
Number of
Utterances
Phase
Signal
Amplitude
Signal
Low-freq
Band
High-freq
Band
No Recurrence no SFA 32 Nerve First Layer Theta Low-Gamma
28 Nerve Nerve Theta Low-Gamma
26 Nerve Nerve Alpha Low-Gamma
21 Nerve First Layer Alpha Low-Gamma
12 Nerve Third Layer Alpha Low-Gamma
Recurrence only 31 Third Layer Third Layer Alpha Low-Gamma
28 Nerve Nerve Theta Low-Gamma
21 Third Layer Third Layer Theta Low-Gamma
20 Nerve Third Layer Alpha Low-Gamma
20 Third Layer Third Layer Beta Low-Gamma
SFA only 51 Third Layer Third Layer Alpha Low-Gamma
42 Second Layer Third Layer Theta High-Gamma
36 Nerve Third Layer Alpha Low-Gamma
33 Second Layer Third Layer Alpha Low-Gamma
26 First Layer Third Layer Theta High-Gamma
Recurrence and SFA 63 Third Layer Third Layer Alpha Low-Gamma
51 Third Layer Third Layer Beta Low-Gamma
27 Third Layer Third Layer Theta Low-Gamma
25 Nerve Nerve Theta Low-Gamma
19 Nerve Third Layer Alpha Low-Gamma
Effects of training and input type on neuronal activity
In order to further understand the emergence of coupled signals, we consider passing speech through an untrained network, as well as passing different types of noise inputs through a trained network. Trained architectures exhibit persisting neuronal activity across layers compared to untrained ones, where the activity almost completely decays after the first layer. In trained networks, noise inputs lead to single neuron firing rate distributions peaking at very low rates and where the activity gradually decreases across layers, as illustrated in Fig. 44. This contrasts with the response to speech inputs seen in Fig. 22 where the activity was sustained across layers with most of the distribution in the gamma range. We tested uniform noise as well as different noise sources (air conditioner, babble, copy machine and typing) from the the MS-SNSD dataset [55]. Compared to a speech input, all noise types yielded reduced average firing rates (from 60 Hz to about 40 Hz) with most of the neurons remaining silent. This highly dynamic processing of information is naturally efficient at attenuating its activity when processing noise or any input that does not induce sufficient synchronisation. Interestingly, babble noises were found in certain cases to induce some significant PAC patterns, whereas other noise types resulted in no coupling at all. Even though babble noises resemble speech and produced some form of synchronisation, they only triggered a few neurons per layer. Overall, we showed that synchronicity of neural oscillations in the form of PAC results from training and is only triggered when passing an appropriate speech input.
[…]
Figure 4: Spiking activity in response to babble noise input. a Input filterbank features and resulting spike trains produced across layers. The model uses a 2 ms time step, 16 CNN channels, 3 layers of size 512, 50% AdLIF neurons, 100% feedforward and 50% recurrent connectivity. b Corresponding single neuron firing rate distributions.
Scaling to a larger dataset
[…]
Training on speech command recognition task
With our experimental setup, the encoder is directly trained to recognise phonemes on TIMIT and subwords on Librispeech. One could therefore assume that the coupled gamma activity emerges from that constraint. In order to test this hypothesis, we run additional experiments on a speech command recognition task where no phoneme or subword recognition is imposed by the training. Instead the model is directly trained to recognise a set of short words. We use the same architecture as on TIMIT, except the average pooling layer is replaced by a readout layer as defined in Bittar and Garner [9] which reduces the temporal dimension altogether, as required by the speech command recognition task. Interestingly, using speech command classes as ground truths still produces significant PAC patterns, especially in the last layer. These results indicate that the emergence of the studied rhythms does not require phoneme-based training and may be naturally emerging from speech processing. Using the second version of the Google Speech Commands data set [57] with 35 classes, we achieve a test set accuracy of 96%, which, to the best of our knowledge, marginally improves upon the current state-of-the-art performance using SNNs of 95.35% [58].
Discussion

### v2

@@ 4.1 Architectural analysis
In order to draw a comparison with the human auditory pathway, we have introduced the physiologically inspired ASR pipeline illustrated in Fig. 2. The proposed hybrid ANN-SNN architecture is trained in an end-to-end fashion on the TIMIT dataset (Garofolo et al., 1993) to predict phoneme sequences from speech waveforms. In the architectural design, we aimed to minimise the complexity of ANN components and favour the central SNN which will be the focus of the oscillation analysis. Here as a preliminary step, we examine how relevant hyperparameters affect the PER. On top of assessing the scalability of our approach to larger networks, we identify the importance of the interplay between recurrence and SFA.
4.1.1 Network scalability
@@
As reported in Table 1, performance improves with added layers, peaking at 4-6 layers before declining, which suggests a significant contribution to the final representation from all layers within this range. Compared to conventional non-spiking RNN encoders used in ASR, our results support the scalability of surrogate gradient SNNs to relatively deep architectures. Additionally, augmenting the number of neurons until about 1,000 per layer consistently yields lower PERs, beyond which performance saturates.
@@
Table 1: Hyperparameter tuning for the number of SNN layers and neurons per layer on the TIMIT dataset. The third column gives both the number of trainable parameters in the multi-layered SNN (left) and in the whole encoder (right). The PERs are reported after 50 training epochs using a 5 ms time step, 16 CNN channels, 50% of AdLIF neurons, 100% feedforward and 50% recurrent connectivity. The performance of the architecture when removing the SNN is also reported (bottom). Bold values indicate the lowest achieved PERs.
@@
4.1.2 Recurrent connections and spike-frequency adaptation
@@
Overall, SFA and recurrent connections individually yield significant error rate reduction, although they respectively grow as $\mathcal{O}(N)$ and $\mathcal{O}(N^{2})$ with the number of neurons $N$. In line with previous studies on speech command recognition tasks (Perez-Nieves et al., 2021; Bittar and Garner, 2022a), our results emphasise the metabolic and computational efficiency gained by harnessing the heterogeneity of adaptive spiking neurons. Furthermore, effectively calibrating the interplay between unit-wise feedback from SFA and layer-wise feedback from recurrent connections appears crucial for achieving optimal performance.
@@
4.1.3 Enforcing Dale’s law
To align with common ANN practice, the previous results were obtained without restricting neurons to be either strictly excitatory or strictly inhibitory. We now train more physiologically inspired models that satisfy Dale’s law, with results presented in Table 3. Although ASR performance decreased (1-4% absolute PER increase), this may simply be due to suboptimal weight initialisation, which is known to affect performance (Li et al., 2024). This could likely be mitigated in future work by using the approach of Rossbroich et al. (2022), who derived fluctuation-driven initialisation schemes compatible with Dale’s law.
Table 3: Ablation experiments for the recurrent connectivity and proportion of neurons with SFA on the TIMIT dataset when additionally using Dale’s law. PERs are reported after 50 epochs using a 2 ms time step, 16 CNN channels, 3 layers, 512 neurons per layer, 100% feedforward connectivity and an excitatory-inhibitory ratio of 1. Bold values indicate the lowest achieved PERs.
Model configuration
Recurrent
connectivity
Proportion of
AdLIF neurons
Number of
parameters
Test PER
Validation PER
No recurrence no SFA 0 0 1.1M / 1.7M 30.7 28.7
Recurrence only 0.5 0 1.5M / 2.1M 23.6 20.6
SFA only 0 0.5 1.1M / 1.7M 25.1 22.9
Recurrence and SFA 0.5 0.5 1.5M / 2.1M 21.2 19.2
4.1.4 Supplementary findings
@@
We observe in Supplementary Table S1 that decreasing the simulation time step does not affect the performance. Although making the simulation of spiking dynamics more realistic, one might have anticipated that backpropagating through more time steps could hinder the training and worsen the performance as observed in standard RNNs often suffering from vanishing or exploding gradients (Bengio et al., 1994). With inputs ranging from 1,000 to over 7,000 steps using 1 ms intervals on TIMIT, our results demonstrate a promising scalability of surrogate gradient SNNs for processing longer sequences.
Secondly, as reported in Supplementary Table S2, we did not observe substantial improvement when increasing the number of auditory nerve fibers past $\sim$5,000, even though there are approximately 30,000 of them in the human auditory system. This could be due to both the absence of a proper model for cochlear and hair cell processing in our pipeline and to the relatively low number of neurons ($<$1,000) in the subsequent layer.
As detailed in Supplementary Table S3, reduced feedforward connectivity in the SNN led to poorer overall performance. This contrasts with our earlier findings on recurrent connectivity, highlighting the distinct functional roles of feedforward and feedback mechanisms in the network.
Additionally, we incorporated trainable delays by replacing the fully connected feedforward matrices with dilated convolutions over the temporal dimension. While including delays resulted in similar overall performance, using the groups parameter to control connectivity led to more parameter-efficient models, as shown in Supplementary Table S4. This method of reducing connectivity proved more effective than our previous approach of randomly masking fully connected matrices.
Finally, Supplementary Table S5 shows our results when using the more popular moving threshold formulation of SFA instead of the AdLIF model. Consistent with our previous findings on speech command recognition (Bittar and Garner, 2022a), the AdLIF implementation of SFA outperforms its moving threshold alternative, with the same number of parameters.
4.2 Oscillation analysis
@@
Based on our previous architectural results that achieved satisfactory speech recognition performance using a physiologically inspired model, we hypothesise that the spiking dynamics of a trained network should, to some extent, replicate those occurring throughout the auditory pathway. Our investigation aims to discern if synchronisation phenomena resembling brain rhythms manifest within the implemented SNNs as they process speech utterances to recognise phonemes.
4.2.1 Synchronised gamma activity produces low-frequency rhythms
@@
As illustrated in Fig. 3 (A), the spike trains produced by passing a test-set speech utterance through the trained architecture exhibit distinct low-frequency rhythmic features in all layers. By looking at the histogram of single neuron firing rates illustrated in Fig. 3 (B), we observe that the distribution predominantly peaks at gamma range, with little to no activity below beta. This reveals that the low-frequency oscillations visible in Fig. 3 (A) actually emerge from the synchronisation of gamma-active neurons. The resulting low-frequency rhythms appear to follow to some degree the intonation and syllabic contours of the input filterbank features and to persist across layers. Compared to the three subsequent layers, higher activity in the auditory nerve comes from the absence of inhibitory SFA and recurrence mechanisms. These initial observations suggest that the representation of higher-order features in the last layer is temporally modulated by lower level features already encoded in the auditory nerve fibers, even though each layer is seen to exhibit distinct rhythmic patterns. In the next section, we focus on measuring this modulation more rigorously via PAC analysis.
Figure 3: Spiking activity in response to speech input. (A) Input filterbank features and resulting spike trains produced across layers. For each layer, the neurons are vertically sorted on the y-axis by increasing average firing rate (top to bottom). The model uses a 2 ms time step, 16 CNN channels, 3 layers of size 512, 50% AdLIF neurons, 100% feedforward and 50% recurrent connectivity with Dale’s law. (B) Corresponding distribution of single neuron firing rates.
4.2.2 Phase-amplitude couplings within and across layers
By aggregating over the relevant spike trains, we compute distinct EEG-like population signals for the auditory nerve fibers and each of the three SNN layers. These are then filtered in the different frequency bands, as illustrated in Fig. 4 (A), which allows us to perform CFC analyses. We measure PAC within-layer and across-layers between all possible combinations of frequency bands and find multiple significant forms of coupling for every utterance.
An example of significant theta low-gamma coupling between the auditory nerve fibers and the last layer is illustrated in Fig. 4 (B). Here input low-frequency modulation is observed to significantly modulate the output gamma activity. This indicates that the network integrates and propagates intonation and syllabic contours across layers through synchronised neural activity along these perceptual cues.
On the majority of utterances, we found significant CFCs between the input waveform and the population signal of the last layer. It is important to note that the synchronisation of neural signals to the auditory envelope emerged without imposing any theta or gamma activity in our network. The optimisation of the PER combined with sufficiently realistic spiking neuronal dynamics therefore represent sufficient conditions to reproduce some broad properties of neural oscillations observed in the brain, suggesting a general functional role of facilitating information processing.
The activity of the final layer of the SNN stands out as the most significantly modulated overall. By architectural construction, modulation in that final layer has the greatest impact on the ASR task, as the spike trains from this layer are converted to phoneme probabilities using a small ANN. A higher number of couplings in the final layer correlates with a decrease in the PER. This suggests that CFCs may be associated with more selective integration of phonetic features, enhanced attentional processes, as well as improved assimilation of contextual information.
Alpha-band oscillations were the most frequently measured, consistent with biological evidence that the alpha rhythm is the most prominent oscillation in spontaneous EEG (Berger, 1929).
More generally, the patterns of coupling between different neural populations and frequency bands were found to differ from one utterance to another. These variations indicate that the neural processing of our network is highly dynamic and depends on the acoustic properties and linguistic content of the input. The observed rich range of intra- and inter-layer couplings suggests that the propagation of low-level input features such as intonation and syllabic rhythm is only one aspect of these synchronised neural dynamics.
Figure 4: Cross-frequency coupling of population aggregated activity. (A) Population signals of auditory nerve fibers (blue) and last layer (orange) filtered in distinct frequency bands. (B) Modulation index and mean vector length metrics as measures of PAC between the theta band of the auditory nerve fibers and the low-gamma band of the last layer.
4.2.3 Impact of Dale’s law on oscillations
As illustrated in Table 4, SNNs that satisfy Dale’s law display significantly higher numbers of CFCs. In biological neural networks, this principle contributes to a more structured and organised form of network dynamics, as each neuron is either excitatory or inhibitory but not both. When applying Dale’s principle to SNNs, it constrains the network with more defined roles for neurons, leading to more coherent oscillatory patterns and more CFCs.
In this study, we use equivalent models for both excitatory and inhibitory neurons, and simply allow them to adapt their parameters within the same fixed ranges during training. We observe that trained values of both membrane and adaptation time constants $\tau_{u}$ and $\tau_{w}$ converge to lower average values across the inhibitory populations compared to the excitatory ones. This suggests that, even with initially equivalent models, the network naturally differentiates the dynamics of excitatory and inhibitory neurons to fulfill their distinct functional roles. Future work could focus on better defining these two types of neurons from the outset, incorporating more biologically plausible initial parameter ranges.
When using Dale’s law, the fixed excitatory-inhibitory (E/I) ratio plays a crucial role in shaping the neuronal dynamics inside the SNN. In our study, we observed that increasing the proportion of inhibitory neurons (E/I = 0.33) resulted in a similar ASR performance ($\sim 1\%$ absolute PER difference) compared to the standard ratio (E/I = 1), but with a reduced average firing rate of 56 Hz instead of 68 Hz for equivalent models with 50% of neurons with SFA and 50% of recurrent connectivity (see Table 4). This finding suggests that a higher ratio of inhibitory neurons can achieve comparable task performance while maintaining a lower overall level of network activity.
Table 4: Effect of Dale’s law, SFA, recurrence and delays on oscillatory activity. Comparison of the oscillatory activity resulting from passing the 64 longest TIMIT test-set utterances through different types of trained networks. The last two columns show the total number of significant PACs summed across all 64 utterances and frequency bands, for intra- and inter-layer relations respectively. All networks use a 2 ms time step, 16 CNN channels, 3 layers, 512 neurons per layer and 100% feedforward connectivity (i.e., groups=1 for delays). The E/I ratio is 1 for models with Dale’s law, except the last one where it is 0.33.
Model type Model configuration
Test PER
$[\boldsymbol{\%}]$
Firing rate
[[[[Hz]]]]
Number of intra-
layer PACs
Number of inter-
layer PACs
AdLIF baseline No recurrence no SFA 26.9 98 133 192
SFA only 23.7 72 213 390
Recurrence only 21.0 76 180 132
Recurrence and SFA 20.5 71 265 177
AdLIF with Dale No recurrence no SFA 30.7 100 513 643
SFA only 25.1 77 564 603
Recurrence only 23.6 67 432 278
Recurrence and SFA 21.2 68 329 290
Recurrence and SFA, E/I=0.33 22.3 56 526 496
AdLIF with delays No recurrence no SFA 26.2 92 116 54
SFA only 23.3 77 91 53
Recurrence only 21.6 68 205 66
Recurrence and SFA 20.8 68 253 60
Moving threshold SFA only 26.3 56 136 131
Recurrence and SFA 22.1 62 182 135
4.2.4 Impact of spike-frequency adaptation and recurrent connections on oscillations
Across all model types (AdLIF or moving threshold SFA, with or without delays and Dale’s law), both SFA and recurrent connections had an overall inhibitory effect, typically reducing the average network firing rate from around 100 Hz to roughly 60 Hz (see Table 4).
This regularisation of the network activity appears to enable more effective parsing and encoding of speech information, as it reliably led to improved ASR performance. While both forms of feedback exhibit an overall inhibitory effect, SFA operates at the individual neuron level whereas recurrent connections act at the layer level.
SFA is known to encourage and stabilise the synchronisation of cortical networks (Crook et al., 1998) and to promote periodic signal propagation (Augustin et al., 2013). In our results, this effect is pronounced in SNNs without Dale’s law, where the inclusion of SFA was consistently associated with a significant increase in the number of measured CFCs. However, when Dale’s law is applied, the overall number of CFCs is significantly higher with no noticeable impact of SFA on CFCs. This suggests that the stricter constraints imposed by Dale’s law may lead to more uniform behaviour in CFCs, thereby reducing the observed influence of SFA.
In models with and without Dale’s law, incorporating recurrent connections was consistently associated with a decrease in the number of inter-layer couplings, indicating more localised synchronisation.
Finally, using the moving threshold formulation of SFA produces a lower firing rate and fewer significant PACs compared to our AdLIF baseline. The lower number of PACs might mean that the moving threshold formulation is less effective than the AdLIF at coordinating these interactions, which could explain the inferior task performance. Nevertheless, the discrepancy between the two approaches could potentially be narrowed with further hyperparameter optimisation.
4.2.5 Impact of delays on oscillations
As illustrated in Table 4, SNNs with delays produced significantly fewer PACs, especially for inter-layer couplings, compared to the baseline with fully connected feedforward matrices. Introducing trainable delays through dilated convolutions allows for temporal dispersion of signals, which may desynchronise neural activity and explain the observed reduction in CFCs.
4.2.6 Effects of training and input type on neuronal activity
In order to further understand the emergence of coupled signals, we consider passing speech through an untrained network, as well as passing different types of noise inputs through a trained network.
Trained architectures exhibit persisting neuronal activity across layers compared to untrained ones, where the activity almost completely decays after the first layer, as illustrated in Figure 5. This decay across layers persists even when increasing the input magnitude up to saturating auditory nerve fibers. This phenomenon can be attributed to the random weights in untrained architectures, which transform structured input patterns into uncorrelated noise, leading to vanishing neuronal activity in deeper layers. Our CFC analysis shows no significant coupling, even in layers with sufficient spiking activity, i.e., within the auditory nerve population and the first layer.
Figure 5: Spiking activity in an untrained network in response to speech input. (A) Input filterbank features and resulting spike trains produced across layers. The model uses a 2 ms time step, 16 CNN channels, 3 layers of size 512, 50% AdLIF neurons, 100% feedforward and 50% recurrent connectivity. (B) Corresponding single neuron firing rate distributions.
In trained networks, noise inputs lead to single neuron firing rate distributions peaking at very low rates and where the activity gradually decreases across layers, as illustrated in Fig. 6 (B). This contrasts with the response to speech inputs seen in Fig. 3 (B) where the activity was sustained across layers with most of the distribution in the gamma range. We tested uniform noise as well as different noise sources (air conditioner, babble, copy machine and typing) from the the MS-SNSD dataset (Reddy et al., 2019). Compared to a speech input, all noise types yielded reduced average firing rates (from 60 Hz to about 40 Hz) with most of the neurons remaining silent. This highly dynamic processing of information is naturally efficient at attenuating its activity when processing noise or any input that does not induce sufficient synchronisation. Interestingly, babble noises were found in certain cases to induce some significant PAC patterns, whereas other noise types resulted in no coupling at all. Even though babble noises resemble speech and produced some form of synchronisation, they only triggered a few neurons per layer (see Fig. 6). Overall, we showed that synchronicity of neural oscillations in the form of PAC results from training and is only triggered when passing an appropriate speech input.
Figure 6: Spiking activity in response to babble noise input. (A) Input filterbank features and resulting spike trains produced across layers. The model uses a 2 ms time step, 16 CNN channels, 3 layers of size 512, 50% AdLIF neurons, 100% feedforward and 50% recurrent connectivity. (B) Corresponding single neuron firing rate distributions.
4.2.7 Scaling to a larger dataset
@@
4.2.8 Training on speech command recognition task
With our experimental setup, the encoder is directly trained to recognise phonemes on TIMIT and subwords on Librispeech. One could therefore assume that the coupled gamma activity emerges from that constraint. In order to test this hypothesis, we run additional experiments on a speech command recognition task where no phoneme or subword recognition is imposed by the training. Instead the model is directly trained to recognise a set of short words. We use the same architecture as on TIMIT, except the average pooling layer is replaced by a readout layer as defined in Bittar and Garner (2022a) which reduces the temporal dimension altogether, as required by the speech command recognition task. Interestingly, using speech command classes as ground truths still produces significant PAC patterns, especially in the last layer. These results indicate that the emergence of the studied rhythms does not require phoneme-based training and may be naturally emerging from speech processing.
Using the second version of the Google Speech Commands data set (Warden, 2018) with 35 classes, we achieve a test set accuracy of 97.05%, which, to the best of our knowledge, improves upon the current state-of-the-art performance using SNNs of 95.35% (Hammouamri et al., 2024).
5 CONCLUSION